# LIBERO-PRO: Towards Robust and Fair Evaluation of Vision-Language-Action Models Beyond Memorization

### v1

@@ 5.3 Does the Model Generalize to New Objects?
To investigate whether the VLA model exhibits true object-level generalization, we conducted tests by altering the manipulated object in the task “pick up the salad dressing and place it in the basket”. As shown in Figure 4, replacing the original salad dressing with an unseen but visually similar variant from outside the training set (second row) did not affect the model’s performance, suggesting a limited degree of generalization to perturbed object instances. However, when substituting salad dressing with an unrelated object (e.g., alphabet soup, third row), the model still executed the same action trajectory, attempting to pick up the new object despite its mismatch with the instruction. This behavior suggests that the model’s apparent generalization is largely driven by rote recall of action trajectories from training rather than genuine grounding of object identity. Moreover, its failure on semantically irrelevant objects highlights the inability to correctly align linguistic references with object semantics. Taken together, these observations indicate that the model’s performance reflects superficial pattern memorization rather than true object-level reasoning.
5.4 Does the Model Generalize to Varied Instructions?
@@
(b) Semantic perturbation: Grab salad dressing and set it in basket.
(c) Meaningless task: fdsgfdsgsd or xxxxxx.
@@
We evaluated whether the VLA model robustly follows natural language instructions. As shown in Figure 5, the model successfully executed the original task (“pick up the salad dressing and place it in the basket”) and also completed the task when the instruction was reformulated with synonymous expressions (“grab the salad dressing and set it in the basket”). Although this outcome may appear to suggest robustness to linguistic variation, when the instruction was replaced with meaningless input (e.g., “xxx”), the model produced the same action sequence, again retrieving and executing the pattern of placing the salad dressing into the basket. This observation indicates that the apparent robustness does not arise from genuine semantic comprehension. Rather, the model relies predominantly on visual context and memorized action trajectories from training, while disregarding the linguistic content of the input. Consequently, the model’s consistent behavior across both paraphrased and nonsensical instructions should not be regarded as successful generalization to semantic perturbations, but instead as evidence of its inability to ground or interpret instructions, revealing an overdependence on rote pattern execution.
5.5 How Sensitive is the Model to Object Placement?
@@
Although all three models achieve around 98% accuracy under the standard LIBERO, our perturbation experiments reveal significant robustness gaps that remain hidden in the original evaluation. As shown in Figure 6, the models are highly sensitive to the initial position of objects: for both OpenVLA and pi0, performance collapses once displacement exceeds 0.2 units, dropping to zero success, while pi0.5—benefiting from a stronger architecture and broader training data—exhibits improved resilience and sustains non-trivial accuracy up to 0.4 units before also collapsing to zero. These results indicate that high scores do not guarantee reliable performance in practice, as even small variations such as placing a cup in a slightly different position can cause complete task failure, underscoring the necessity of LIBERO-PRO to systematically reveal such robustness gaps and provide a more faithful comparison across models.
5.6 Is the Model Truly Understanding Tasks?
@@
To assess whether the model demonstrates genuine task comprehension or merely memorizes action trajectories, we designed a sequence of three progressively structured tasks (as shown in Figure 7). In the first task, the agent is instructed to pick up the bowl and place it on the stove, and the model executes this sequence successfully, indicating its ability to reproduce simple manipulation behaviors. In the second task, the instruction is to turn on the stove, and again the model succeeds, suggesting competence in carrying out isolated atomic actions. However, in the third task, which composes these two operations into a longer sequence—first picking up the bowl, then placing it on the stove, and finally turning the stove on—the model consistently fails. Rather than demonstrating adaptive reasoning over the composed instruction, the model replays spurious trajectories resembling those seen during training, leading to unsuccessful execution. This discrepancy reveals a critical limitation: the model does not truly understand either the semantic goal of the task or the causal relationships between subtasks, but instead exhibits rote reproduction of memorized patterns from the training distribution, and the failure on even modestly composed instructions underscores the fragility of current approaches and highlights the need for evaluation protocols that probe beyond single-step or template-like behaviors.

### v2

@@ 2.1 Visual perceptual hallucinations
Executing the grasp on an unrelated object. As shown in Figure 2(b), when we replace the target salad dressing bottle with an unrelated item—alphabet soup—the model still executes the instruction “Pick up the salad dressing and place it in the basket” by reaching toward and grasping the alphabet soup. This behavior indicates that the model does not reliably bind the instruction to the correct visual entity. Instead, it appears to infer the intended action from the overall scene context and subsequently replay a familiar trajectory, regardless of whether the visual evidence supports the action.
Executing the grasp despite the object’s removal. A similar error emerges when the target object is entirely removed from the scene. As illustrated in Figure 2(c), even though the salad dressing is absent, the model still performs the grasping motion toward the location where the object typically appears. This “grasping the air” behavior further indicates that the policy does not perform online visual verification of object existence before initiating the action.
2.2 Instruction understand hallucinations
Ignoring meaningless instructions.
@@ 2.2 Instruction understand hallucinations
(b) Meaningless task: fdsgfdsgsd or xxxxxx.
@@
Figure 3(b), when we replace the original instruction with a meaningless character sequence (e.g., “fdsgfdsgsd” or “xxx”), the model still executes nearly the same action trajectory as in the original task “Pick up the salad dressing and place it in the basket.” Rather than recognizing the absence of valid semantic content, the model proceeds to grasp the salad dressing and attempt the placement action without any hesitation. This behavior indicates that the policy is largely insensitive to linguistic semantics and that the instruction input contributes little to the decision-making process. Instead, the model appears to rely primarily on visual context to retrieve a memorized visuomotor pattern, further reinforcing that its apparent instruction-following capability does not stem from genuine language grounding.
2.3 Location perception hallucinations
@@
Our qualitative results show that the policy is highly sensitive to the object’s initial position: when the manipulated object is displaced away from the coordinates commonly observed in the training demonstrations, the model consistently fails to execute the grasp. Even small offsets lead the end-effector to move toward the original, memorized location rather than the object’s actual position. This behavior indicates that the policy does not perform online spatial reasoning or object localization.
As shown in Figure 4, the models are extremely sensitive to shifts in the object’s initial position: for both OpenVLA and pi0, task success collapses once the displacement exceeds 0.2 units, dropping sharply to zero thereafter. Even the stronger pi0.5 model—benefiting from a larger architecture and more diverse training data—only maintains non-trivial performance up to roughly 0.4 units before also degrading to zero.
2.4 Lack of task understanding
@@
To evaluate whether the model possesses genuine task-level understanding rather than merely reproducing isolated demonstrations, we design a sequence of three progressively structured tasks (Figure 5). In the first task—pick up the bowl and place it on the stove—the model succeeds, indicating it can replicate a simple manipulation trajectory. In the second task—turn on the stove—the model again performs reliably, suggesting competence with single-step atomic actions.
However, when the two subtasks are combined into a single multi-step instruction—pick up the bowl, place it on the stove, and then turn on the stove—the model consistently fails. Rather than assembling the subtasks into a coherent execution plan, it replays partial or spurious trajectories reminiscent of the training distribution. This discrepancy reveals that the model lacks compositional reasoning: it does not understand the temporal or causal structure linking the subtasks, but instead depends on memorized action fragments. Even such modest task compositions exceed its generalization capacity, highlighting the fragility of current VLA models.
These observations suggest that the strong in-distribution performance of current Visual-Language-Action (VLA) models may rely heavily on superficial regularities in the training data. Although a VLA policy is intended to learn a mapping $\pi:(\mathcal{O},\mathcal{L})\rightarrow\mathcal{A}$ from visual observations and language instructions to actions, our results indicate that the learned policy often fails to robustly ground actions in the actual observation-instruction pair. Instead, it appears to exploit shortcuts such as canonical scene layouts, recurring object appearances, typical object positions, and repeated motion trajectories. This provides a plausible explanation for the observed failures: the model grasps absent or incorrect objects, ignores meaningless instructions, reaches toward memorized locations under positional perturbations, and fails to compose individually learned subtasks into coherent multi-step behaviors. Overall, these results suggest that current VLA models may still depend more on shortcut feature-action correlations than on robust and compositional visuomotor understanding.
3 Problem Definition
3.1 Formal VLA Task Definition
We formalize a vision-language-action (VLA) task instance as
$\mathcal{T}=(l,e,G),$ (1)
where $l\in\mathcal{L}$ is a natural-language instruction, $e\in\mathcal{E}$ denotes an environment instance, $\mathcal{E}$ is the space of environment instances, and $G:\mathcal{S}\times\mathcal{L}\to\{0,1\}$ is a binary success predicate indicating whether a state satisfies the task specified by the instruction. An environment instance is defined as
$e=(\mathcal{W},\mathcal{O},R_{0}),$ (2)
where $\mathcal{W}$ denotes the visual environment context (e.g., background, textures, and lighting conditions), $\mathcal{O}$ denotes the set of interactive objects in the scene, and $R_{0}$ denotes the initial spatial configuration of the objects. Let $s_{0}$ denote an initial state consistent with $R_{0}$. A policy $\pi_{\theta}(a_{t}\mid o_{t},l)$ induces trajectories according to
$a_{t}\sim\pi_{\theta}(\cdot\mid o_{t},l),\qquad s_{t+1}\sim P(\cdot\mid s_{t},a_{t}),$ (3)
where $o_{t}$ denotes the agent observation at time $t$, and $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ denotes the environment dynamics. We evaluate a policy by its success probability within a finite horizon $H$:
$J(\pi_{\theta};\mathcal{T})=\mathbb{P}_{\pi_{\theta},\,e}\!\left[\exists t\in\{0,\dots,H\},\;G(s_{t},l)=1\right],$ (4)
where the probability is taken over trajectories induced by $\pi_{\theta}$ under the environment $e$. This objective captures the joint requirements of language grounding, perception, and action.
3.2 Current LIBERO Evaluation Protocol
While LIBERO has been an influential benchmark for VLA research, its current evaluation protocol exhibits limited distributional shift between training and test tasks. Under our formulation, each task is specified by an instruction $l$, an environment instance $e=(\mathcal{W},\mathcal{O},R_{0})$, and a success predicate $G$.
In LIBERO, a test task $\mathcal{T}^{\prime}\in\mathcal{T}_{\mathrm{test}}$ is typically constructed from a corresponding training task $\mathcal{T}\in\mathcal{T}_{\mathrm{train}}$ such that
$l^{\prime}=l,\qquad\mathcal{W}^{\prime}=\mathcal{W},\qquad\mathcal{O}^{\prime}=\mathcal{O},\qquad G^{\prime}=G,$ (5)
with variation arising only from small perturbations to the initial spatial configuration $R_{0}$, i.e., $R_{0}^{\prime}$ differs from $R_{0}$ only by minor perturbations.
Consequently, test tasks remain close to training tasks, and evaluation primarily reflects robustness to small variations in object configurations rather than generalization to novel task compositions or environments. As a result, strong performance may largely arise from interpolation within a narrow neighborhood of the training distribution, potentially overestimating a model’s ability to generalize under broader distribution shifts.
4 Methodology
4.1 Proposed Perturbation-Based Evaluation Framework
Figure 6: Overview of LIBERO-PRO. LIBERO-PRO extends LIBERO with five perturbation types: object, configuration, instruction, task, and environment. These perturbations vary object attributes, object placements, language descriptions, task goals/object sets, and visual context, respectively, enabling systematic evaluation under diverse distribution shifts. To ensure feasibility, task perturbations are excluded from cross-type combinations, while the other perturbation types can be freely composed.
To enable more principled evaluation of VLA models, we design a perturbation-based framework that introduces controlled variations across four orthogonal dimensions: object attributes, initial spatial configurations, task instructions, and environments. Given a task $\mathcal{T}=(l,e,G)$ with environment instance $e=(\mathcal{W},\mathcal{O},R_{0})$, we generate perturbed variants $\mathcal{T}^{(k)}=\phi_{k}(\mathcal{T})$, where $k\in\{\mathrm{O},\mathrm{R},\mathrm{L},\mathrm{E}\}$.
Object attribute perturbation ($k=\mathrm{O}$) modifies non-essential visual attributes of objects, such as color, texture, or size, while preserving semantic equivalence. The resulting task is
$\mathcal{T}^{(\mathrm{O})}=(l,e^{(\mathrm{O})},G),\qquad e^{(\mathrm{O})}=(\mathcal{W},\mathcal{O}^{(\mathrm{O})},R_{0}).$
Initial configuration perturbation ($k=\mathrm{R}$) alters the absolute and relative positions of objects while maintaining physical plausibility. The perturbed task is
$\mathcal{T}^{(\mathrm{R})}=(l,e^{(\mathrm{R})},G),\qquad e^{(\mathrm{R})}=(\mathcal{W},\mathcal{O},R_{0}^{(\mathrm{R})}).$
Instruction perturbation ($k=\mathrm{L}$) includes both semantic- and task-level variations. Semantic perturbations rephrase the instruction while preserving task intent:
$\mathcal{T}^{(\mathrm{L}_{\mathrm{sem}})}=(l^{(\mathrm{L})},e,G).$
Task perturbation ($k=\mathrm{T}$) modifies the target object or action while remaining within the training distribution. The perturbed task is
$\mathcal{T}^{(\mathrm{T})}=(l,e,G^{(\mathrm{T})}).$
Environment perturbation ($k=\mathrm{E}$) modifies the visual environment context, such as background, lighting, or texture, without altering task feasibility. The resulting task is
$\mathcal{T}^{(\mathrm{E})}=(l,e^{(\mathrm{E})},G),\qquad e^{(\mathrm{E})}=(\mathcal{W}^{(\mathrm{E})},\mathcal{O},R_{0}).$
4.2 Perturbation Constraints
To ensure that each perturbed task remains semantically coherent and executable, we impose two classes of constraints. First, perturbations are bounded within a dimension-specific neighborhood:
$x^{(k)}\in\mathcal{N}_{k}(x),\qquad\mathrm{dist}_{k}(x,x^{(k)})\leq\delta_{k},$
where $x$ denotes a task component, $x^{(k)}$ its perturbed counterpart, and $\delta_{k}$ controls the perturbation magnitude along dimension $k$. Second, we enforce non-trivial task-level separation:
$d_{\mathrm{TV}}(\mathcal{T},\mathcal{T}^{(k)})>\epsilon,$
where $d_{\mathrm{TV}}$ denotes the total variation distance, ensuring that perturbations induce meaningful behavioral changes. By varying $\delta_{k}$, we generate multiple perturbed variants for each category and characterize policy sensitivity to perturbation intensity.
4.3 LIBERO-PRO Evaluation Benchmark
Building upon the above framework, we introduce LIBERO-PRO, a plug-and-play benchmark for systematic evaluation across multiple dimensions of variation, including object attributes, spatial configurations, instructions, tasks, and environments, as well as their combinations.
For object variation, we modify object color, texture, and scale. For configuration variation, we alter object placements while preserving physical validity. For instruction variation, we generate paraphrased task descriptions. For task variation, we redesign tasks to introduce new object sets and goals while maintaining executability. For environment variation, we vary visual context such as backgrounds and lighting.
To avoid infeasible combinations, task-level perturbations are excluded from cross-type compositions, while the remaining perturbation types may be freely combined to construct diverse evaluation settings.
5 Experiments
5.1 Settings
Figure 7: Main results on LIBERO-PRO. Despite high average success on standard LIBERO tasks, current VLA models fail sharply under controlled perturbations, particularly position and task shifts. Robustness to object, semantic instruction, and environment changes is limited and varies across models and task suites.
We conduct experiments on LIBERO-PRO, which applies controlled perturbations from four perspectives: object properties, initial positions, task instructions (covering both semantic and structural variations), and environments. We evaluate three representative models—OpenVLA[8], pi0[1], and pi0.5[6]. For fairness, OpenVLA is assessed using four official checkpoints trained on different LIBERO suites, whereas pi0 and pi0.5 are evaluated with a single official checkpoint. Consistent with the original LIBERO protocol, we set the number of evaluation episodes to 50 per task.
5.2 Main Results
We conducted a comprehensive evaluation of state-of-the-art vision-language-action (VLA) models [1, 8] using the LIBERO-PRO evaluation suite, which introduces perturbations across four dimensions: objects, goals, spatial configurations, and task sets (see Tables 7). Our findings are as follows: (1) Failure under simple perturbations. Despite achieving success rates above 90% on the standard LIBERO benchmark, models nearly collapse under changes to object positions or minor task modifications, even when constructed from training components. (2) Illusory robustness. Apparent stability against object and instruction perturbations is largely due to overfitting rather than genuine task understanding. (3) Model-specific sensitivity. Environmental perturbations affect models differently, depending on the degree of environmental divergence and reliance on memorized contexts. (4) Subtle capability differences. While most models fail under perturbations, Pi0.5 achieves a 0.38 success rate in the libero-goal task under position changes, compared to 0 for OpenVLA and Pi0, revealing differences masked by standard benchmark scores. Overall, these results show that current VLA models, despite excelling on the standard LIBERO, lack true robustness and generalization—highlighting the need for rigorous evaluation frameworks.
6 Related Work
Vision-Language-Action Models. Vision-Language-Action (VLA) models have rapidly emerged as a central paradigm for embodied AI, aiming to unify perception, natural language understanding, and action execution within a single framework. Recent years have witnessed an explosion of approaches, ranging from large-scale pre-trained models that leverage multimodal corpora to task-specific architectures tailored for robotic manipulation and navigation[14, 10, 2]. While these efforts have significantly advanced the field, they also underscore the urgent need for standardized evaluation benchmarks[13, 15, 7]. Without such benchmarks, it becomes difficult to conduct fair and reproducible comparisons across methods, hindering the cumulative progress of VLA research.
Benchmarks for VLA Evaluation. A number of benchmarks have been developed to enable systematic evaluation of VLA models, including RLBench [7], CALVIN [15], RoboCasa [16], and BridgeData [19]. These benchmarks provide valuable testbeds for task execution under vision and language guidance, yet their adoption has remained relatively limited within the broader community. In contrast, LIBERO [13] has quickly emerged as the dominant benchmark and is now the most widely adopted protocol for evaluating VLA models. Its standardized task suite and unified reporting metrics have made it the de facto standard for performance comparison, to the extent that virtually all recent VLA studies report results on LIBERO. As such, LIBERO not only serves as the “common currency” of evaluation in this field, but also exerts a strong influence on research directions and claims of progress.
Limitations of Current Evaluation Practices. Several recent studies have noted that the high scores reported on LIBERO often fail to translate into reliable performance, even on tasks with only minor variations. This has raised serious concerns about the validity of its evaluation protocol[21, 3, 11]. In response, some researchers have explored stronger algorithms or proposed stress tests [4, 5, 18, 9, 20]to encourage more robust VLA systems. However, the field still lacks a systematic and standardized framework for robust evaluation under LIBERO. The majority of works continue to rely on the original, flawed setup, reporting numbers that may not reflect genuine capability. To address this gap, we propose LIBERO-PRO, a plug-and-play extension of LIBERO designed to foster fair comparison across models and to advance research on building VLA systems that can withstand reasonable and practically necessary perturbations.